# Energy consumption of code small language models serving with runtime engines and execution providers

### v1

@@ 7 Limitations & threats to validity
Our study has various limitations, primarily due to its focus on open-source SLMs, specific runtime engines, and execution providers. The findings may not generalize to other model types or serving configurations. Although the provided replication package allows for other models and configurations to be easily added, the conclusions are limited to the examined setup and code SLMs context. Our findings may not generalize to all deep learning systems, but they offer a foundation for optimizing and analyzing energy-efficient serving configurations. These results are specific to the particular used hardware, described in Section 4.6, and future work could mitigate this by incorporating a wider range of hardware, including different GPUs and CPUs, to enhance the robustness of the findings.
@@
Additionally, this study focused on the code generation task using prompts derived from the HumanEval dataset [23]. While HumanEval is one of the most widely used evaluation datasets, its scope may not fully capture the diversity of real-world scenarios. This limitation is particularly relevant given that our research work and development of serving infrastructure were initiated in late 2023. Expanding the scope to include more comprehensive and newer benchmarks, such as SWE-bench dataset [76], is a clear future work option to potentially provide broader insights into code SLMs serving configurations.
@@
Lastly, the reliance on the Transformers library for implementation introduces some limitations in flexibility, as other libraries might yield different performance results. Future work could explore using alternatives to assess generalizability across different implementations.
@@
Conclusion validity. The statistical power of our experiments was carefully designed to ensure the reliability of our findings. To address this, we employed statistical methods, adhered closely to a methodically planned experimental procedure as detailed in section 4, and ensured reproducibility through comprehensive documentation of our methods and data in the replication package. Considering various factors, treatments, and runs, we utilized a total of 600 (12 code SLMs, 5 deep learning serving configurations, 10 runs) samples to respond to our research questions, each sample representing a set of 164 inference requests. This approach, coupled with the execution of ten experimental runs and ensuring minimal background software interference, aims to fortify the reliability of our results. The procedure for replicating our work is thoroughly described, with datasets and the replication package openly accessible for validation.
@@
Construct validity. To mitigate mono-method bias, we utilized three different profiling tools and a wattmeter to cross-verify the energy consumption recorded by the CPU (via energibridge) and GPU (via nvidia-smi) against the total energy reported by the wattmeter, ensuring the validity of our energy measurements. Additionally, we included twelve state-of-the-art SLMs in our study. Concerning model correctness, we did not evaluate this aspect as we employed well-established, previously validated models. All experimental outputs are provided in the replication package for comprehensive examination.
@@
Internal validity. Potential internal validity threats such as hardware condition changes over successive experimental runs were addressed by implementing a five-minute period between each configuration and experiment run. This protocol helps maintain consistent hardware conditions across all tests. To minimize risks associated with the implementation of the inference process, we used default configurations provided by HuggingFace.
@@
External validity. Our study’s external validity is potentially limited by the specific subset of deep learning serving configurations and models used. The models and configurations were selected based on their current relevance and availability, but they may not encompass all possible variations in deep learning serving scenarios. To mitigate this, we incorporated the two most commonly used execution providers available to us and four well-known runtime engines.
@@ 8 Conclusions and future work
In this work, we studied deep learning serving configurations, as duplets consisting of a runtime engine and an execution provider. We evaluated the energy consumption, execution time, and computing-resource utilization across five distinct configurations using twelve different code SLMs. Our observations reveal that:
@@
    TorchScript engine shows promise as an optimal choice for specific scenarios, though it may require further enhancements to fulfill its potential for code generation.
@@
We anticipate that future advancements in optimized runtime engines will enhance these metrics and address existing limitations, including error-proneness, to make them even more robust and efficient.
@@
Our research highlights distinct performance advantages among the various configurations. Specifically, the CUDA execution provider paired with TORCH consistently yielded the best overall performance. Conversely, when utilizing the CPU execution provider, both OV and ONNX runtime engines enhanced performance metrics significantly. In configurations with the CUDA execution provider and TORCH, energy consumption using this configuration can be from 10.84% up to 62.01%, and execution times from 10.26% up to 52.16% comparing to worst CUDA execution provider runtime engine (ONNX runtime). For setups using the CPU execution provider with OV or ONNX, we observed the energy consumption can be from 27.96% up to 91.02%, and execution times from 28.79% up to 79.04% compared to the baseline TORCH engine. Notably, the ONNX runtime engine did not exhibit superior performance compared to the baseline TORCH configuration within CUDA execution provider configurations.
@@
Future research should replicate these experiments across a diverse range of deep learning tasks, extending beyond computer vision to encompass various natural language processing tasks like summarization and text classification. Such studies will not only confirm the generalizability of our findings but may also reveal insights specific to different tasks. Additionally, we recommend investigating newly developed runtime engines and execution providers. As technology advances, these newer configurations are likely to enhance performance and energy efficiency, further promoting sustainable deep learning practices.

### v2

@@ 7 Limitations and threats to validity
Our study has various limitations, primarily due to its focus on open-source SLMs, specific runtime engines, and execution providers. The findings may not generalize to other model types or serving configurations. Although the provided replication package allows for other models and configurations to be easily added, the conclusions are limited to the examined setup and code generation SLMs context. Our findings may not generalize to all deep learning systems, but they offer a foundation for optimizing and analyzing resource utilization-efficient serving configurations. These results are specific to the particular used hardware, described in Section 4.6, and future work could mitigate this by incorporating a wider range of hardware, including different GPUs and CPUs, to enhance the robustness of the findings.
@@
Additionally, this study focused on the code generation task using prompts derived from the HumanEval dataset [25]. While HumanEval is one of the most widely used evaluation datasets, its scope may not fully capture the diversity of real-world scenarios. For instance, prompts in these type of scenarios can span multiple files, integrate external libraries, or involve different releases. This limitation is particularly relevant given that our research work and development of serving infrastructure were initiated in late 2023. Expanding the scope to include more comprehensive and newer benchmarks, such as SWE-bench dataset [68], is a clear future work option to provide broader insights into code generation SLMs serving configurations.
@@
Lastly, the reliance on the Transformers library for implementation introduces some limitations in flexibility, as other libraries might yield different performance results. For instance, ONNX’s used version still lacks kernels for operators in CUDA execution provider and falls back to the CPU execution provider in the version we used [96], causing data transfers and addition of new nodes to the executed model. This imbalance can overstate resource utilization advantages for $\langle TORCH,CUDA\rangle$ within CUDA configurations. Future work could explore using alternatives to assess generalizability across different implementations.
@@
Conclusion validity. The statistical power of our experiments was carefully designed to ensure the reliability of our findings. To address this, we employed statistical methods, adhered closely to a methodically planned experimental procedure as detailed in Section 4, and ensured reproducibility through comprehensive documentation of our methods and data in the replication package. Considering various factors, treatments, and runs, we utilized a total of 600 (12 code generation SLMs, 5 deep learning serving configurations, 10 runs) samples to respond to our research questions, each sample representing a set of 164 inference requests. This approach, coupled with the execution of ten experimental runs and ensuring minimal background software interference, aims to fortify the reliability of our results. The procedure for replicating our work is thoroughly described, with datasets and the replication package openly accessible for validation.
@@
Construct validity. To mitigate mono-method bias, we utilized two different profiling tools and a wattmeter to cross-verify the energy consumption recorded by the CPU (via energibridge) and GPU (via nvidia-smi) against the total energy reported by the wattmeter, ensuring the validity of our energy measurements. Additionally, we included twelve state-of-the-art SLMs in our study. Concerning model correctness, we did not evaluate this aspect as we employed well-established, previously validated models. All experimental outputs are provided in the replication package for comprehensive examination.
@@
Internal validity. Potential internal validity threats such as hardware condition changes over successive experimental runs were addressed by implementing a five-minute period between each configuration and experiment run. This protocol helps maintain consistent hardware conditions across all tests. To minimize risks associated with the implementation of the inference process, we used default configurations provided by HuggingFace. While the repetition strategy and idle-time control minimized variability and helped ensure fair comparisons, fixed-order execution may still introduce minor risks related to system temperature changes or background activity fluctuations. Future studies could include shuffling of configurations to further mitigate such risks.
@@
External validity. Our study’s external validity is potentially limited by the specific subset of deep learning serving configurations and models used. The models and configurations were selected based on their current relevance and availability, but they may not encompass all possible variations in deep learning serving scenarios. To mitigate this, we incorporated the two most commonly used execution providers available to us and four well-known runtime engines. Although JIT was part of the initial design, it was excluded from evaluation due to the error-prone inference results.
Moreover, external validity is influenced by the representativeness of the used dataset itself. HumanEval contains basic to intermediate coding challenges, and our input dataset of 10 to 15 tokens captures only the short prompt auto-complete code task. Consequently, our findings are suited to interactive auto-complete scenarios and may not extend to generation that requires longer context, the synthesis of complete functions, or multi-file conversations. Future work should therefore adopt larger prompts and benchmarks that span multiple files to measure serving resource utilization under those conditions.
@@ 8 Conclusions and future work
The role of infrastructure-aware evaluation is expected to become increasingly important. For high-risk systems (as classified in the EU AI Act), the legislation introduces provisions that may require providers to meet energy-efficiency standards and to document the energy consumption of AI systems throughout their lifecycle, including during the inference phase (Art. 40) [97]. These obligations are likely to encourage more detailed and standardized reporting practices. Future research can support this shift by developing methodologies to isolate not only the energy impact but also the resource utilization impact of serving infrastructure components, to clarify the distribution of resource consumption across lifecycle phases (e.g., training versus inference), and to analyze the software stack involved in inference. Figure 1 provides a general visualization of the deep learning serving stack, which includes components such as runtime engines and execution providers, highlighted in red. These efforts can also contribute to the definition of benchmarks aligned with regulatory expectations. This regulatory direction may foster the adoption of more transparent, auditable, and resource-efficient deployment practices, creating new opportunities to enhance the sustainability of deep learning systems at scale.
In this work, we studied deep learning serving configurations, as duplets consisting of a runtime engine and an execution provider. We evaluated the energy consumption, execution time, and computing-resource utilization across five distinct configurations using twelve different code generation SLMs. Our observations reveal that:
@@
    JIT engine shows promise as an optimal choice for specific scenarios, though it may require further enhancements to fulfill its potential for code generation.
@@
The observed dominance of the $\langle TORCH,CUDA\rangle$ configuration reflects its maturity, extensive support, and integration within the PyTorch and HuggingFace ecosystem, rather than inherent model-specific optimization. Nonetheless, optimization effectiveness might vary for other models and tasks. To support serving infrastructure decisions in real-world tasks, we provide context-aware guidelines that help practitioners prioritize among energy consumption, execution time, and computing-resource utilization depending on system-level constraints. These guidelines serve as a starting point for deploying similar decoder-based SLMs in various scenarios. For example, execution time may be prioritized in latency-critical settings such as IDE-based auto-complete code task, while energy efficiency becomes more relevant in battery-constrained or cost-sensitive deployments. In computing-resource utilization limited systems, minimizing RAM or VRAM usage may justify sacrificing some speed or energy performance. By framing serving infrastructure selection as a trade-off decision, our guidance enables engineers to tailor serving configurations to operational goals. We anticipate that future advancements in optimized runtime engines will enhance these metrics and address existing limitations, including error-proneness, to make them even more robust and efficient. Consequently, our study should be viewed as a best-case snapshot for $\langle TORCH,CUDA\rangle$ on used hardware. As kernel coverage converges and libraries configurations such as IOBinding are tuned, alternative engines are likely to approach, or even surpass, this baseline [80].
@@
Our research highlights distinct resource utilization advantages among the various configurations. Specifically, the CUDA execution provider paired with TORCH consistently yielded the best overall resource utilization. Conversely, when utilizing the CPU execution provider, both OV and ONNX runtime engines enhanced resource utilization metrics significantly. In configurations with the CUDA execution provider and TORCH, energy consumption using this configuration can be from 10.84% up to 62.01%, and execution times from 10.26% up to 52.16% comparing to worst CUDA execution provider runtime engine (ONNX runtime). For setups using the CPU execution provider with OV or ONNX, we observed the energy consumption can be from 27.96% up to 91.02%, and execution times from 28.79% up to 79.04% compared to the baseline TORCH engine. Notably, the ONNX runtime engine did not exhibit superior resource utilization compared to the baseline TORCH configuration within CUDA execution provider configurations.
While JIT can offer performance or resource utilization benefits, it struggles with code generation. Future efforts might investigate alternative just-in-time approaches capable of handling branching logic more effectively. Overcoming these limitations would help to serve a broader range of use cases involving diverse inputs and execution paths.
Our findings explicitly underscore the complex, nuanced relationships among energy consumption, execution time, and computing resource utilization in SLMs inference infrastructures. We highlighted clear scenarios where optimizing execution time did not necessarily optimize energy consumption, reinforcing the explicit need for direct energy monitoring in practical settings. Subsequent studies should explicitly focus on investigating these trade-offs across broader deep-learning scenarios and model architectures, aiming to provide more generalized frameworks and actionable guidelines for infrastructure optimization.
This work reinforces the importance of treating infrastructure components as architectural elements in deep learning system design. We recommend integrating the resource utilization profiling of serving infrastructure directly into development workflows and MLOps pipelines, helping software teams continuously evaluate trade-offs between energy, time, and computing-resource utilization across system updates and deployments.
@@
Future research should replicate our experiments across a broader spectrum of deep learning tasks and input lengths. Because the present study concentrated on short prompts that represent auto-complete code task, further work is needed to verify whether the configurations advantages we observed hold when generating large-context or conversational code generation. Benchmarks such as SWE bench [68] and realistic multiple file tasks can reveal how serving configurations scale in these richer settings. Beyond code generation, investigations in areas like summarization and text classification will confirm the generalizability of our findings and may uncover insights that are specific to particular tasks. Additionally, we recommend investigating newly developed runtime engines and execution providers. As technology advances, these newer configurations are likely to enhance performance and resource utilization efficiency, further promoting sustainable deep learning practices.